OpenStudy (anonymous):

What equation models this problem? Colin divided his baseball cards in 6 equal piles.He puts the cards from five into albums.How many cards are in his collection if he put 150 baseball cards into albums

OpenStudy (anonymous):

@steve816

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

25

OpenStudy (anonymous):

@liljohn10 What equation model

OpenStudy (anonymous):

dividing

OpenStudy (anonymous):

@liljohn10 like 140/6b=5 140/5b=6 5/6b=140 6/5b=140

OpenStudy (anonymous):

@liljohn10 Which one

OpenStudy (anonymous):

@steve816 @liljohn10

OpenStudy (anonymous):

ask @steve816

OpenStudy (anonymous):

He left the question!

OpenStudy (anonymous):

not me

OpenStudy (anonymous):

I mint to